OpenStudy (krystal.bright21):

The half-life of a substance is how long it takes for half of the substance to decay or become harmless (for certain radioactive materials). The half-life of a substance is 123 days and there is an amount equal to 150 grams now. What is the expression for the amount A(t) that remains after t days, and what is the amount of the substance remaining (rounded to the nearest tenth) after 365 days?

myininaya (myininaya):

You know we are going to have an exponential equation. \[A(t)=A_0 e^{kt}\] where A_0 is the present amount of substance and the is time in days

myininaya (myininaya):

we are given when t=123 days we have A(t) is half the present amount of substance

OpenStudy (krystal.bright21):

A(t) = 150(0.5)t/123, 19.2 grams remaining A(t) = 150(0.5)123t, 1.4 grams remaining A(t) = 123(150)(0.5)t, 0.0 gram remaining A(t) = 150(0.5)123/t, 118.8 grams remaining

myininaya (myininaya):

\[\frac{A_0}{2}=A_0 e^{k 123}\] solve for k

OpenStudy (krystal.bright21):

I'm new to this.. I don't know how

myininaya (myininaya):

you should see you can divide both sides by A_0 and write \[\frac{1}{2}=e^{ 123 k}\] can you solve this for k?